# The 1/3 Geometric Constant: Scale Invariance and the Origin of ’Missing Energy’ in 3D Quantum Fragmentation

### v4

@@ V.1 Empirical Validation: The $H_{2}^{+}$ Energy Discrepancy
Figure 4: Energy distribution analysis and empirical benchmarks. (a) Proton KER spectra for $H_{2}^{+}$ dissociative ionization. Numerical results represent 1D (dashed lines, Ref. Yue and Madsen (2013)) and 3D (solid lines, this work) computations. Experimental data are denoted by triangle markers (Ref. Pavičić et al. (2005)) and dot markers (Ref. Rudenko et al. (2005), after applying a logarithmic transformation). Black curves correspond to ground-state ($n=1$) contributions; red curves include the summation of populated excited states. (b) Energy distribution $P(E)$ as a function of principal quantum number $n$ ($n=1,2,3$) for fixed $Q=1,\zeta=1$. (c) Comparison of the historical 210Bi beta-decay spectrum (triangle markers, Ref. Ellis and Wooster (1927)) with the theoretical $P(E)$ distribution at $\alpha=0.16$ (black dashed line).
@@
The physical validity of the 0.33 geometric landmark is most strikingly evidenced by the long-standing energy discrepancy in attosecond $H_{2}^{+}$ fragmentation. As shown in the Fig. 4(a), traditional 1D semiclassical models Yue and Madsen (2013) predict a KER peak between 10101313 eV, significantly overestimating experimental observations. In contrast, our 3D simulations (solid lines) yield a primary peak near $3$ eV, in excellent agreement with experimental calorimetry Pavičić et al. (2005); Rudenko et al. (2005).
@@
Our model replicates the observed $\approx 7$ eV shift in $H_{2}^{+}$ fragmentation without requiring external field effects or multi-photon dynamics, identifying the $4\pi r^{2}$ volume element as the primary driver of the discrepancy. While the large mass of the escaping protons ($M_{p}\approx 1836$ a.u.) would typically drive the system toward a classical energy ratio, their initial localization within the $1\sigma_{g}$ molecular manifold provides a critical symmetry-induced compensation. Because the protons are initially constrained along the internuclear axis—a spatial distribution templated by the axial $\psi\propto r\cos^{m}\theta e^{-\zeta r}$ symmetry—the angular components of the Laplacian ($\nabla^{2}\psi\propto m/r^{2}$) significantly enhance the local "quantum pressure." This effectively neutralizes the mass scaling in our dimensionless coupling $\alpha=MQ/\zeta$. This mechanism maintains the heavy fragment system near the $\alpha\approx 1$ geometric equilibrium, where the volume weighting sequesters the majority of the potential energy. Consequently, the detected proton energy represents only one-third of the integrated mean ($\langle E\rangle\approx 11$ eV), confirming that the "missing energy" in molecular calorimetry is a direct manifestation of 3D quantum geometry.
This interplay between symmetry and mass is further illuminated when compared to subatomic regimes. While axial alignment in molecules provides a "quantum pressure" that counteracts fragment mass, nuclear fragmentation generally lacks this directional constraint. In the spherically symmetric environment of the nucleus, the large fragment mass becomes the dominant driver of the scaling constant $\alpha\to\infty$, shifting the detection ratio toward the sharp-peaked classical limit observed in Alpha decay.
V.2 Nodal Structures and the Mechanism of Spectral Broadening
@@
Beyond the ground state ($n=1$), higher principal quantum numbers introduce additional radial nodes that fundamentally alter the topology of the energy distribution $P(E)$. As illustrated in the Upper Left panel of Fig. 4(b), simulations for $n=1,2,$ and $3$ (at fixed $Q=1,\zeta=1,M=1$) reveal that while the position of the primary low-energy peak remains relatively stable, increasing $n$ triggers the emergence of secondary features at higher energies. Physically, the local kinetic energy $T_{\text{local}}(r)$ incorporates terms related to the curvature of these additional radial nodes, redistributing probability density into regions of higher local kinetic energy and creating a multi-modal distribution. In the high-$n$ limit, this secondary feature can undergo a dominance shift, eventually surpassing the primary peak as the most prominent characteristic of the spectrum. We note that the varying high-energy truncation points in these numerical curves are artifacts of grid-spacing limits during the discretization of increasingly complex orbital geometries.
@@
This nodal redistribution provides a definitive geometric explanation for the spectral broadening observed in high-intensity $H_{2}^{+}$ experiments in Fig. 4(a). When excited states ($n>1$) are populated, the 3D model identifies secondary high-energy features between 4455 eV (red solid curve), which are entirely absent in 1D approximations. These features serve as a direct signature of the radial nodal structures in the initial-state wavepackets. As shown in Fig. 4(a), the inclusion of these excited-state contributions explains the characteristic broadening and structured high-energy tails observed in experimental data (black and red dot markers, Ref. Rudenko et al. (2005)). Consequently, while the $1/3$ landmark defines the primary detection peak, the population of higher shells ($\zeta_{\text{eff}}$) provides the underlying geometric mechanism for spectral width evolution and the observed calorimetric structure of the fragmentation.
V.3 The Geometric Origin of Missing Energy in Subatomic Physics
@@
The identification of the $M=\alpha\zeta/Q$ scaling law provides a provocative new lens for interpreting energy anomalies in nuclear physics. Historically, the continuous electron spectrum observed in beta decay—characterized by an average energy $\langle E\rangle\approx 1/3E_{\text{max}}$—was the primary catalyst for the neutrino hypothesis Ellis and Wooster (1927). Notably, the morphology of the beta spectrum (Fig. 4(c)) is remarkably well-replicated by our model in the low-to-mid energy range ($0.5E_{\text{max}}$) using a coupling constant of $\alpha=0.16$. This value corresponds to a high-localization ($\zeta$) regime, consistent with the high atomic number ($Z$).
The divergence of the fit at higher energies is expected, as this region approaches the relativistic regime where the non-relativistic Laplacian operator used in our current $P(E)$ derivation overestimates the energy drive. Furthermore, the nuclear Q-value imposes a hard phase-space cutoff at the endpoint energy that truncates the theoretical high-energy tail. Despite these specific high-energy corrections, our results demonstrate that the $R_{E}=0.33$ ratio represents a fundamental geometric baseline for energy detection in localized 3D systems. This universal constraint offers a unified explanation for perceived energy suppression across diverse physical scales, from molecular Coulomb explosions to the quenching factors observed in dark matter direct detection experiments Levin and others (1991); Aprile and others (XENON Collaboration) (2023)
V.4 Scaling Limits: Comparing Alpha and Beta Decay
@@
The universal nature of the scaling law is further underscored by the contrasting spectral morphologies of Alpha Krane (1987) and Beta decay. While the low-mass regime of Beta decay (210Bi) resides near the $\alpha\approx 1$ landmark—where 3D geometric masking is maximal—the emission of heavy helium nuclei in Alpha decay represents the high-$\alpha$ limit of our framework. As the scaling constant $\alpha=MQ/\zeta$ increases due to the massive fragment $M$, our model predicts a geometric sharpening of $P(E)$ and a convergence of $R_{E}$ toward the classical limit. This explains why Alpha spectra appear as nearly mono-energetic lines with high detection efficiency, in contrast to the broad, “missing energy” distributions of Beta decay. While Alpha transitions involve complex $Q$-value dynamics beyond our current potential-inversion framework, the shift from masked to visible energy peaks confirms that the observability of subatomic transitions is fundamentally governed by the dimensionless scaling of the 3D volume element.
V.5 Unified Mechanism of Spectral Broadening and Orbital Contraction
The spectral distribution $P(E)$ is critically governed by the localization parameter $\zeta$. As illustrated in our parametric study (Fig. 2), an increase in $\zeta$ induces a rapid broadening of the high-energy tail. In high-$Z$ nuclei, the proximity of inner-shell electrons effectively scales the $\zeta$ parameter, leading to the spectral broadening observed in traditional nuclear studies Wilkinson (1990).
@@ V.5 Unified Mechanism of Spectral Broadening and Orbital Contraction
For superheavy nuclei ($Z>100$), the extreme localization (high $\zeta$) increases the probability of detecting ultra-high energy fragments. This phenomenon aligns with experimental observations of relativistic orbital contraction Henning (2012). While currently interpreted via Fermi theory Pershina (2016), our model suggests these relativistic shifts are consistent with the geometric pressure inherent to high-$\zeta$ localized states, providing a scale-invariant link between atomic and relativistic nuclear dynamics.
VI Conclusion
@@
In this work, we have demonstrated that the persistent discrepancy between theoretical energy budgets and experimental peaks in quantum fragmentation is not a technical limitation, but a fundamental topological artifact of three-dimensional space. By synthesizing localized Slater-type orbitals with full-dimensional TDSE benchmarks, we have established that the $4\pi r^{2}$ volume element acts as an inescapable geometric filter. This mechanism inherently masks a significant portion of a system’s interaction energy, imposing a universal peak-to-mean bound of $R_{E}<0.5$.
@@
The discovery of the invariant scaling law, $\alpha=(MQ)/\zeta$, provides a unified framework for energy reconstruction across twelve orders of magnitude. We have shown that the 7 eV "missing energy" in attosecond $H_{2}^{+}$ fragmentation is a direct manifestation of this geometric constraint, resolving the anomaly without the need for external field corrections. Most significantly, the alignment between this 0.33 landmark and the historical energy morphology of beta decay suggests that the 1/3 ratio is a universal baseline for 3D detection.
@@
Our findings provide a transformative perspective on the interpretation of “missing energy” in subatomic physics. By mapping different decay modes onto our universal master curve, we show that the discrepancy in Beta decay and the high-resolution efficiency of Alpha decay are two limits of the same geometric scaling. While the neutrino remains a cornerstone of the Standard Model, our results reveal that the baseline for energy detection in localized states is inherently suppressed by the topology of the initial state. Because localization ($\zeta$) typically increases with particle mass ($M$), subatomic systems are naturally driven toward regimes where $R_{E}$ is highly sensitive to 3D masking. Consequently, geometric masking must be rigorously decoupled from signatures of invisible flux. This work shifts the focus of future calorimetry from the search for undetected mass to the calibration of the intrinsic geometric filters of our three-dimensional quantum world.
@@ Appendix A Methods
A.1 Bohmian Local Energy Derivation
@@
A.2 Numerical Validation against TDSE
@@
A.3 Symmetry-Dependent Local Kinetic Energy
@@
A.4 Derivation of the Universal Scaling Law
@@ 3. Homogeneous Invariance
A.5 $H_{2}^{+}$ Simulation Protocols and Parameter Selection

### v5

@@ IV Discussion: From Molecular Calorimetry to Subatomic Decay
Figure 4: Energy distribution analysis and empirical benchmarks. (a) Proton KER spectra for $H_{2}^{+}$ dissociative ionization. Numerical results represent 1D (dashed lines, Ref. Yue and Madsen (2013)) and 3D (solid lines, this work) computations. Experimental data are denoted by triangle markers (Ref. Pavičić et al. (2005)) and dot markers (Ref. Rudenko et al. (2005), after applying a logarithmic transformation). Black curves correspond to ground-state ($n=1$) contributions; red curves include the summation of populated excited states. (b) Energy distribution $P(E)$ as a function of principal quantum number $n$ ($n=1,2,3$) for fixed $Q=1,\zeta=1$. (c) Comparison of the historical 210Bi beta-decay spectrum (triangle markers, Ref. Ellis and Wooster (1927)) with the theoretical $P(E)$ distribution at $\alpha=0.16$ (red dotted line).
IV.1 Empirical Validation: Resolving the $H_{2}^{+}$ Energy Anomaly
@@
The physical validity of the 0.33 geometric landmark is most strikingly evidenced by the long-standing energy discrepancy in attosecond $H_{2}^{+}$ fragmentation. As shown in Fig. 4(a), traditional 1D semiclassical models Yue and Madsen (2013) predict a kinetic energy release (KER) peak between 10101313 eV (black dashed line), significantly overestimating experimental observations (black triangles). In contrast, our 3D formulation yields a primary peak near 33 eV (black solid line), in excellent agreement with experimental calorimetry (black triangles) Pavičić et al. (2005); Rudenko et al. (2005).
@@
Our model replicates the observed $\approx 7$ eV shift without invoking external field corrections or complex multi-photon dynamics, identifying the $4\pi r^{2}$ volume element as the primary driver of the discrepancy. While the large mass of the escaping protons ($M_{p}\approx 1836$ a.u.) would classically drive the system toward a high $R_{E}$ ratio, their initial localization within the $1\sigma_{g}$ molecular manifold provides a critical symmetry-induced compensation. Because the protons are constrained along the internuclear axis—a distribution templated by the axial $\psi\propto r\cos^{m}\theta e^{-\zeta r}$ symmetry—the angular components of the Laplacian ($\nabla^{2}\psi\propto m/r^{2}$) significantly enhance the local “quantum pressure.” This effect effectively neutralizes the mass scaling in the dimensionless coupling $\alpha=MQ/\zeta$, maintaining the heavy fragment system near the $\alpha\approx 1$ geometric equilibrium. Consequently, the detected proton energy represents only one-third of the integrated mean ($\langle E\rangle\approx 11$ eV), confirming that the “missing energy” in molecular calorimetry is a direct manifestation of 3D quantum topology.
IV.2 Nodal Structures and Spectral Broadening
@@
Beyond the ground state ($n=1$), higher principal quantum numbers introduce radial nodes that alter the energy distribution $P(E)$. As illustrated in Fig. 4(b), simulations for $n=1,2,$ and 33 reveal that while the primary low-energy peak remains stable, increasing $n$ triggers secondary high-energy features. Physically, the local kinetic energy $T_{\text{local}}(r)$ incorporates terms related to the curvature of these additional nodes, redistributing probability density into higher kinetic energy regimes and creating a multi-modal spectrum.
@@
This nodal redistribution provides a geometric explanation for the spectral broadening observed in high-intensity $H_{2}^{+}$ experiments (red and black dots in Fig. 4(a)). When excited states ($n>1$) are populated, our 3D model identifies secondary features between 4455 eV (red solid curve in Fig. 4(a)) which are absent in 1D approximations. These features serve as a direct signature of the radial nodal structures in the initial-state wavepacket. Thus, while the $1/3$ landmark defines the primary detection peak, the population of higher-order shells provides the underlying geometric mechanism for spectral width evolution.
IV.3 The Cascading Geometric Mask: A Two-Process Model of $\beta$ Decay
@@
The identification of the $\alpha=MQ/\zeta$ scaling law provides a provocative lens for interpreting energy anomalies in nuclear physics. Historically, the continuous electron spectrum in $\beta$ decay—characterized by a peak and average energy significantly below the endpoint ($\langle E\rangle\approx 1/3E_{\text{max}}$)—was the catalyst for the neutrino hypothesis Ellis and Wooster (1927). While our model replicates the spectral morphology using a coupling of $\alpha=0.16$ in the non-relativistic regime(Fig. 4(c)), we propose that the total energy deficit arises from a cascading geometric mask involving two distinct quantum processes.
In this framework, we consider the fragmentation as a sequential interaction between a primary nuclear fragment (particle AA) and a secondary inner-shell electron (particle $B$). We characterize the event through the following hierarchy:
  1. 1.
    The Geometric Gate (Process 1): Particle AA is born from a highly localized nuclear environment. Before interacting with the atomic shell, its own energy distribution is subject to the 3D geometric baseline. We assume that the effective interaction energy AA provides to the system is constrained by its own spectral peak ($E_{\text{peak}}\approx 1/3Q$, marked by red trigangles in Fig. 4(c)). This pre-filtering defines the maximum energy budget available for secondary ejections.
  2. 2.
    Potential Inversion and Overlap (Process 2): As particle AA echelons toward the atomic shell, particle $B$ experiences a sudden transition from an attractive to a repulsive potential ($V\to+Q$). The transfer of energy from AA to $B$ is governed by the quantum overlap integral $\langle\psi_{B}|\hat{V}|\psi_{A}\rangle$.
Crucially, the interaction is dominated by regions of maximum spatial density. Because both AA and $B$ are 3D wavefunctions filtered by the $4\pi r^{2}$ volume element, their maximum overlap occurs precisely at their respective radial peaks. This implies a **Quantum Selection Rule**: only the peak-energy component of particle AA’s wavepacket can effectively couple with particle $B$.
The remaining two-thirds of the total nuclear energy, sequestered in the high-curvature radial tails of the wavefunctions, remains topologically inaccessible to the secondary electron. Consequently, the detected $\beta$ spectrum is a convolution of two geometric constraints. This explains why the observed energy is tethered to the 0.33 landmark not as a result of weak-force kinematics, but as a structural necessity of 3D wave-coupling. By mapping these decay modes onto our universal master curve, we demonstrate that the "missing energy" in localized subatomic systems is a direct manifestation of this cascading geometric sequestration.
IV.4 Scaling Limits: Comparing Alpha and Beta Decay
@@
The universal nature of the scaling law is further underscored by the contrasting morphologies of $\alpha$ and $\beta$ decay. While the low-mass regime of $\beta$ decay resides near the $\alpha\approx 1$ landmark—where geometric masking is maximal—the emission of heavy helium nuclei in $\alpha$ decay represents the high-$\alpha$ limit. As $M$ increases, our model predicts a geometric sharpening of $P(E)$ and a convergence of $R_{E}$ toward the classical limit. This explains why $\alpha$ spectra appear as nearly mono-energetic lines with high detection efficiency, in contrast to the broad distributions of $\beta$ decay. The shift from masked to visible energy peaks confirms that the observability of subatomic transitions is fundamentally governed by the dimensionless scaling of the 3D volume element.
IV.5 Orbital Contraction and Relativistic Scaling
@@
The spectral distribution $P(E)$ is critically governed by the localization parameter $\zeta$. In high-$Z$ nuclei, the proximity of inner-shell electrons effectively scales $\zeta$, leading to the spectral broadening observed in traditional nuclear studies Wilkinson (1990). For superheavy nuclei ($Z>100$), extreme localization increases the probability of detecting ultra-high energy fragments, a phenomenon consistent with experimental observations of relativistic orbital contraction Henning (2012). While currently interpreted via Fermi theory, our model suggests these shifts are consistent with the geometric pressure inherent to high-$\zeta$ localized states, providing a scale-invariant link between atomic and relativistic nuclear dynamics.
V Conclusion
@@
In this work, we have demonstrated that the persistent discrepancy between theoretical energy budgets and experimental peaks in quantum fragmentation is not a consequence of technical limitations, but a fundamental topological artifact of three-dimensional space. By synthesizing localized Slater-type orbitals with full-dimensional TDSE benchmarks, we have established that the $4\pi r^{2}$ volume element acts as an inescapable geometric filter. This mechanism inherently masks a significant portion of a system’s interaction energy, imposing a universal peak-to-mean bound of $R_{E}<0.5$ on all 3D fragmentation spectra.
@@
The discovery of the invariant scaling law, $\alpha=MQ/\zeta$, provides a unified framework for energy reconstruction across twelve orders of magnitude. We have shown that the 7 eV “missing energy” in attosecond $H_{2}^{+}$ fragmentation is a direct manifestation of this geometric constraint, resolving a long-standing anomaly without the need for external field corrections or complex dynamical modeling. Most significantly, the alignment between the 0.33 landmark and the historical energy morphology of $\beta$ decay suggests that the $1/3$ ratio is a universal baseline for energy detection in localized 3D systems.
@@
Our findings provide a transformative perspective on the interpretation of energy anomalies in subatomic physics. By mapping disparate decay modes onto a universal master curve, we demonstrate that the broad spectrum of $\beta$ decay and the high-resolution efficiency of $\alpha$ decay represent two limits of the same geometric scaling. While the neutrino remains a fundamental component of the Standard Model, our results reveal that the baseline for energy detection in localized states is inherently suppressed by the topology of the initial-state wavefunction.
Because localization ($\zeta$) correlates with particle mass ($M$), subatomic systems are naturally driven toward regimes where the detection ratio $R_{E}$ is highly sensitive to 3D masking. Consequently, these geometric effects must be rigorously decoupled from signatures of invisible flux or new particles. This work shifts the focus of future calorimetry from the search for undetected mass to the precision calibration of the intrinsic geometric filters inherent to our three-dimensional quantum world.
@@ Appendix A Methods
A.1 Sudden Approximation and Hamiltonian Transmutation
The applicability of our 3D fragmentation model to high-energy decay processes is justified via the sudden approximation. In attosecond molecular physics, the rapid ionization of bonding electrons transforms an attractive potential into a repulsive landscape before the nuclear wavefunctions can adiabatically adjust.
This framework can be extended to the nuclear regime, where reaction energies are sufficiently high to ensure that emitted particles reach ultra-relativistic velocities before the surrounding electronic environment can respond. In $\beta$-decay, for instance, the transmutation timescale is significantly shorter than the characteristic orbital period of the inner-shell electrons. Consequently, the electronic wavefunctions remain spatially invariant (or "frozen") during the event. The newly born high-speed electron introduces a localized Coulombic repulsion that suddenly competes with the nuclear attraction. This process, analogous to the Migdal "shake-off" or Auger-like transitions, forces a reconfiguration of the electronic cloud.
A.2 Bohmian Local Energy Derivation
@@
A.3 Numerical Validation against TDSE
@@
A.4 Symmetry-Dependent Local Kinetic Energy
@@
A.5 Derivation of the Universal Scaling Law
@@ 3. Homogeneous Invariance
A.6 $H_{2}^{+}$ Simulation Protocols and Parameter Selection